Assume that the reaction represented by the equation A → products is second order in A with a rate constant of 8.5×10-2 M-1·s-1. What is the concentration of A after a reaction time of 5s if the initial concentration of A was 7.0 M? Enter your result in M, without the unit, to a precision of 0.1M.
Answer #1

(1/CA)-(1/7M)=8.5*10-2M-1s-1*5s

(1/CA)=(8.5*10-2M-1*5)+(1/7)M-1

(1/CA)=0.56785714286M-1

CA=1.761